# Free-Space Optical Channel Turbulence Prediction: A Machine Learning Approach

### v1

@@ VI Conclusion
In conclusion, this research presented an empirically-validated, novel approach for predicting discrete turbulence levels of an optical data channel utilizing machine learning, specifically the XGBoost model. The method consistently achieved an accuracy of $>98$% so long as the dynamic changes in turbulence were allowed to stabilize over about 1 minute. A temporal analysis revealed that prediction accuracy could actually degrade with more training data during turbulence stabilization periods. The study also provides valuable insights into optimizing training strategies by examining the impact of the number of bits per data instance and the number of data instances on prediction accuracy. Looking ahead, future research directions include exploring machine learning models that better adapt to real-time dynamics of atmospheric turbulence and perhaps provide more quantitative data about the channel. Tradeoff studies considering the hardware constraints for practical ML deployment may also be compared to existing methodologies. By bridging the gap between machine learning and FSO communication, this research contributes to the development of more resilient and efficient communication systems.

### v2

@@ VI Conclusion
This research presented a novel, empirically-validated approach to affirm the possibility of predicting discrete turbulence levels within an optical data channel by utilizing ML, specifically the XGBoost model. The method consistently achieved an accuracy of $>98$% so long as the dynamic changes in turbulence were allowed to stabilize over about 1 min. A temporal analysis revealed that prediction accuracy could actually degrade with more training data during turbulence stabilization periods. The study also provides valuable insights into optimizing training strategies by examining the impact of the number of bits per data instance and the number of data instances on prediction accuracy. Looking ahead, future research directions include exploring ML models that better adapt to real-time dynamics of turbulence and provide more quantitative data about the channel. Tradeoff studies considering the hardware constraints for practical ML deployment and training, may also be compared to existing methodologies. By bridging the gap between ML and FSO communication, this research contributes to the development of more resilient and efficient communication systems.